### v2

# Long Context Modeling with Ranked Memory-Augmented Retrieval

Abstract
Effective long-term memory management is crucial for language models handling extended contexts. We introduce the Enhanced Ranked Memory Augmented Retrieval (ERMAR) framework, which dynamically ranks memory entries based on relevance. Unlike prior models, ERMAR employs a novel relevance scoring mechanism and a pointwise re-ranking model for key-value embeddings, inspired by learning-to-rank techniques in information retrieval. By integrating historical usage patterns and adaptive retrieval, ERMAR achieves state-of-the-art results on standard benchmarks, demonstrating superior scalability and performance in long-context tasks.
Long Context Modeling with Ranked Memory-Augmented Retrieval
Ghadir Alselwi1, Hao Xue1, Shoaib Jameel2, Basem Suleiman1, Hakim Hacid3, Flora D. Salim1, Imran Razzak4,1 1University of New South Wales, Sydney, NSW, Austalia 2University of Southampton, Southampton, UK 3Technology Innovation Institute, Abu Dhabi, UAE 4Mohamed Bin Zayed University of Artificial Intelligence, Abu Dhabi, UAE {g.alselwi,hao.xue1,b.suleiman, flora.salim}@unsw.edu.au m.s.jameel@southampton.ac.uk, hakim.hacid@tii.ae, imran.razzak@mbzuai.ac.ae
1 Introduction
Large Language Models (LLMs) face a fundamental limitation in processing long-context scenarios due to the quadratic complexity of attention mechanisms and increasing memory demands during generation Vaswani (2017); Tworkowski et al. (2024). Consider a scenario in an automated customer service system: A customer reports an issue with their printer, referencing a setup process from a previous conversation that occurred two hours ago. After 50 messages of troubleshooting, the customer mentions that the same error from the beginning has resurfaced. Traditional LLMs, constrained by their context window, would struggle to access the crucial earlier context about the initial setup process, leading to inconsistent or incomplete responses, Figure 1. It is well known that handling extended contexts remains a significant challenge, particularly in applications requiring document analysis and sustained dialogue interactions.
The recent MemLong Liu et al. (2024) architecture stores and accesses historical context through basic chunk-level memory operations. The memory bank model is a large, non-trainable store of past context representations. Instead of re-computing representations for all past tokens every time, these representations are pre-computed and stored. Given the current context, MemLong retrieves relevant segments from the memory bank. It uses a dot product similarity search to find the memory entries most related to the current context. This allows the model to focus only on the most pertinent past information, rather than processing the entire history. However, its treatment of all key-value (K-V) pairs with equal weight, regardless of their contextual relevance, often leads to information overload and reduced retrieval precision. This limitation becomes particularly evident in scenarios requiring context management.
We have developed a novel model that addresses the aforementioned limitations by building upon Memlong Liu et al. (2024), a publicly available baseline on GitHub [^1] https://github.com/Bui1dMySea/MemLong. Our Enhanced Ranked Memory Augmented Retrieval (ERMAR) model has a novel relevance scoring mechanism that fundamentally improves context retrieval and utilization for K-V embeddings. Unlike MemLong, ERMAR employs multiplication Cao et al. (2007) to compute relevance scores, enabling a more nuanced and context-aware assessment of semantic alignment between queries and stored memory. ERMAR also incorporates a re-ranking mechanism that dynamically reorders K-V embeddings based on their relevance scores, ensuring that the most pertinent information is prioritized during retrieval. This re-ranking process, combined with an adaptive retrieval system that integrates historical usage patterns, allows ERMAR to capture subtle contextual relationships better and refine memory prioritization. As shown in Figure 1, ERMAR processes incoming queries and long-context conversations through a novel ranking architecture, employing K-V pairs ranking ( $K_{0}$-$V_{0}$, $K_{1}$-$V_{1}$,…,$K_{i}$-$V_{i}$ ) and corresponding embeddings to perform semantic search and ranking of relevant historical information.
Figure 1: Our novel proposed ERMAR system. Note the difference from the MemLong architecture where we have introduced a novel Reranking model.
Our novel ERMAR model introduces three key improvements: (i) A semantic similarity metric to measure contextual alignment between query embeddings and key-value pairs; (ii) A weighted scoring function that considers content similarity and contextual relevance; and (iii) Integration of historical usage patterns to refine relevance assessment.
2 Related Work
The challenge of enabling language models to effectively process extended contexts has driven research across multiple domains. We organize related work into key areas that inform our ERMAR framework.
2.1 Memory-Augmented Neural Networks
Early neural memory architectures like Neural Turing Machines Graves et al. (2014) and Differentiable Neural Computers Graves et al. (2016) established the principle of external memory mechanisms beyond parameter storage. Recent memory-augmented architectures have focused on language modeling tasks. RetroMAE Xiao et al. (2022) demonstrated promising results but struggled with semantic coherence when retrieved passages lacked contextual relevance. Memorizing Transformers Wu et al. (2022) introduced dedicated memory tokens for cross-sequence information storage, showing improvements on long-range dependency tasks but facing scalability limitations. The recent MemLong architecture Liu et al. (2024) represents a significant advance by storing historical context through chunk-level memory operations using a non-trainable memory bank. However, its uniform treatment of all key-value pairs regardless of contextual relevance often leads to information overload and reduced retrieval precision, particularly in scenarios requiring nuanced context management.
2.2 Retrieval-Augmented Generation and Long-Context Modeling
Retrieval-augmented generation (RAG)Lewis et al. (2020) pioneered integrating dense passage retrieval with generative models, while Fusion-in-Decoder (FiD)Izacard and Grave (2020) improved efficiency through independent passage processing. REALM Guu et al. (2020) introduced end-to-end learning of retrieval and generation, and DPR Karpukhin et al. (2020) established dense passage retrieval standards. However, these methods typically focus on static corpora rather than dynamic context-aware memory management. The quadratic complexity of attention mechanisms has driven research into efficient long-context architectures. Sparse attention mechanisms such as Longformer Beltagy et al. (2020) and BigBird Zaheer et al. (2020) reduce computational complexity while maintaining model capabilities through selective attention patterns. Position encoding adaptations like RoPE Su et al. (2024) and ALiBi Press et al. (2022) have enhanced models’ ability to handle longer sequences. YARN Peng et al. (2023) further advanced this through dynamic position embeddings, demonstrating reliable generalization up to 128k tokens.
2.3 Learning-to-Rank and Information Retrieval
Our ERMAR framework draws inspiration from learning-to-rank techniques in information retrieval. Traditional ranking approaches like BM25 struggle with semantic similarity, leading to neural ranking models using learned representations. Pointwise ranking approaches Cao et al. (2007) predict relevance scores for query-document pairs, directly inspiring our relevance scoring mechanism. Dense retrieval methods like DPR Karpukhin et al. (2020) and ColBERT Khattab and Zaharia (2020) demonstrate the effectiveness of learned dense representations for ranking through similarity scores—a paradigm we adapt for memory entry ranking. BERT-based re-ranking models Nogueira and Cho (2019) show that sophisticated re-ranking significantly improves retrieval quality. This two-stage retrieve-then-rerank paradigm directly influences our design, where we first retrieve candidate memory entries then apply learned re-ranking.
2.4 Current Limitations and Gaps
Despite significant progress, current approaches face limitations that motivate ERMAR:
Static Memory Management: Most approaches use fixed memory structures that don’t adapt to content importance or usage patterns, leading to inefficient memory utilization.
Uniform Memory Treatment: Existing methods treat all memory entries equally, lacking mechanisms to prioritize more relevant information.
Limited Semantic Understanding: Memory systems often rely on simple similarity metrics without considering contextual nuance or temporal relevance.
Scalability Trade-offs: Current approaches face difficult trade-offs between memory capacity, retrieval accuracy, and computational efficiency.
Our ERMAR framework addresses these limitations through dynamic relevance scoring, adaptive memory management, and sophisticated re-ranking mechanisms inspired by information retrieval techniques, providing a more principled approach to long-context memory management.
3 Our Novel ERMAR Model
Figure 2 illustrates our ERMAR framework and Figure 1 presents the contextual ranking mechanism of key, value pairs components that enable effective retrieval. ERMAR maintains consistency through frozen lower layers and selective parameter updating. ERMAR: (i) stores important information from earlier parts of the text; (ii) assigns relevance scores to stored information based on its importance to the current context, and (iii) retrieves only the most relevant historical information when needed. Our relevance scoring is analogous to attention, allowing the model to focus on important parts of the memory. There is also a “loose” pointwise connection because the primary objective is sequence likelihood.
Figure 2: The architecture diagram for ERMAR.
Let $\mathcal{V}$ be a finite vocabulary, and
$\mathbf{x}=(x_{1},\ldots,x_{n})\in\mathcal{V}^{n}$
a token sequence with preceding context $\mathbf{x}_{<i}$. The embedding function
$\mathcal{E}:\mathcal{V}^{*}\rightarrow\mathbb{R}^{d_{\text{ret}}}$
maps sequences to retrieval space. We introduce a memory function $\mathcal{M}$ augmented with a relevance scoring mechanism:
$\mathcal{M}:\underbrace{\mathbb{R}^{d_{\text{model}}}}_{\text{keys}}\times\underbrace{\mathbb{R}^{d_{\text{model}}}}_{\text{values}}\times\underbrace{\mathbb{R}^{d_{\text{ret}}}}_{\text{embeddings}}\rightarrow\mathcal{S}$
Relevance Score: Given a query embedding $\mathbf{q}\in\mathbb{R}^{d_{\text{ret}}}$ and a matrix of key embeddings
$\mathbf{K}=[k_{1},\ldots,k_{m}]\in\mathbb{R}^{m\times d_{\text{ret}}}$
(where each row $k_{i}\in\mathbb{R}^{d_{\text{ret}}}$ corresponds to a key embedding), the relevance score is:
$\alpha(\mathbf{q},\mathbf{K})=\text{softmax}\left(\frac{\mathbf{q}\mathbf{K}^{\top}}{\sqrt{d_{\text{ret}}}}\right)$ (1)
Here, $\sqrt{d_{\text{ret}}}$ normalizes the similarity scores to prevent excessively large values. The relevance score $\alpha(\mathbf{q},\mathbf{K})$ can be interpreted as a probability distribution over the keys, where each entry $\alpha_{i}$ represents the relative importance (or attention weight) of the $i$-th key to the query $\mathbf{q}$. This score is used to rank memory entries based on their importance to the current query.
Ranked Key-Value Pairs: Each embedding $\mathbf{e}$ maintains a ranked set of key-value pairs:
$\mathcal{R}_{\text{ranked}}(\mathbf{e})=\{(K_{j},V_{j},s_{j})\}_{j=1}^{m}$
where $s_{j}=\alpha(\mathbf{e},K_{j})$ is the relevance score between the embedding $\mathbf{e}$ and the key $K_{j}$.
ERMAR objective function is formulated as:
Given a sequence $\mathbf{x}$, maximize:
$\mathcal{L}(\theta)=\sum_{i=1}^{n}p_{\theta}(x_{i}\mid\mathcal{R}_{\text{RSAR}}(t_{i},s),\mathbf{x}_{<i})$
subject to:
$s_{i}$ $=\mathcal{M}(K_{1:i-1},V_{1:i-1};\mathcal{E}(t_{1:i-1}))$
$t_{i}$ $=\text{text}(c_{\lceil i/\tau\rceil})$
$\alpha_{i}$ $=\alpha(\mathcal{E}(t_{i}),K_{1:i-1})$
where $\alpha_{i}$ guides key-value pair selection, and $p_{\theta}$ represents the model’s probability distribution.
For new content $(K_{\text{n}},V_{\text{n}})$, update the memory state as:
$s_{i+1}=\begin{cases}\mathcal{M}(K_{\text{n}},V_{\text{n}};\mathcal{E}(t_{\text{n}}))&\text{if }|s_{i}|<\text{capacity},\\ M_{\text{u}}(s_{i},K_{\text{n}},V_{\text{n}},\alpha_{\text{n}})&\text{otherwise},\end{cases}$
where $M_{\text{u}}$ prunes the least relevant entries based on historical scores, specifically by ranking the scores and pruning those with the lowest values relative to the current context.
We now develop the Relevance Scoring with Adaptive Retrieval (RSAR). This approach dynamically ranks memory entries based on their importance to the current query, significantly improving the retrieval process. The relevance score, $\alpha(\mathbf{q},\mathbf{K})$, as defined in equation 1, is used to rank memory entries.
RSAR enhances the memory module by introducing ranked key-value entries, represented as $(K_{j},V_{j},s_{j})$, where $s_{j}$ denotes the relevance score for each entry. These scores enable the system to prioritize the most relevant information during retrieval while maintaining computational efficiency. To ensure optimal memory utilization, a pruning strategy is applied to remove less relevant entries. Specifically, entries with scores below a predefined threshold are discarded, preserving only the most critical context.
The enhanced retrieval mechanism is expressed as:
$R_{\text{RSAR}}(t_{q},s)=\text{TopK}\left\{\text{sim}(E(t_{q}),e)\cdot\max_{j}s_{j}\mid e\in s\right\}$
where $E(t_{q})$ represents the encoded query, and the operation identifies the top-$K$ relevant entries based on their scores. This mechanism efficiently retrieves the most relevant information, even for extended contexts.
3.1 Experimental Setup
3.1.1 Datasets
We fine-tuned ERMAR on the SlimPajama dataset Fu et al. (2024), a high-quality, deduplicated corpus designed for long-context tasks. It contains 84.7K training rows, making it a compact yet effective resource for pre-training and fine-tuning. The dataset was preprocessed with a sliding window approach using 512-token strides to ensure comprehensive coverage of long sequences.
Performance evaluation was conducted on three benchmark datasets: WikiText-103 Merity et al. (2016) (4,358 test rows), PG-19 Rae et al. (2019) (100 test rows), and Proof-Pile Azerbayev et al. (2023) (46.3K test rows). Performance was measured across context lengths from 1024 to 32768 tokens, using perplexity on the last 2048 tokens Yen et al. (2024) following standard evaluation protocols.
3.1.2 Base Model Architecture
We fine-tuned OpenLLaMA-3B, a pre-trained LLM with rotational position encoding Su et al. (2024), using LoRA Hu et al. (2021) for parameter-efficient fine-tuning. The model architecture consists of $L=26$ transformer layers, $H=32$ attention heads, and hidden dimension $d=3200$ (note: the original $d=100$ appears to be a typo). The 13th layer serves as the memory layer where historical context is stored, while layers [14, 18, 22, 26] are augmented with retrieval mechanisms to access stored memories.
3.1.3 Model Configuration
We fine-tuned OpenLLaMA-3B using LoRA Hu et al. (2021) for parameter-efficient training. The model uses layer 13 as the memory layer for storing historical context, while layers [14, 18, 22, 26] are augmented with retrieval mechanisms. ERMAR employs a memory capacity of 32,768 key-value pairs with BGE-M3 embeddings for semantic similarity computation. Complete training hyperparameters and configuration details are provided in AppendixAppendix.
3.1.4 Baseline Models
ERMAR was evaluated against state-of-the-art models across two parameter scales to ensure comprehensive comparison. The 7B models include LLaMA-2-7B Touvron et al. (2023b) as a standard transformer baseline, LongLoRA-7B-32k Chen et al. (2023) which employs sparse attention mechanisms for 32k-token contexts, and YARN-128k-7B Peng et al. (2023) featuring dynamic position embeddings that support up to 128k tokens.
For the 3B parameter scale, we compared against OpenLLaMA-3B Touvron et al. (2023a) as the base architecture, LongLLaMA-3B Tworkowski et al. (2024) evaluated in two retrieval configurations (4 and 18 memory entries), MemLong-3B Liu et al. (2024) as our direct baseline with chunk-level memory operations, and Phi3-128k Abdin et al. (2024) which demonstrates strong performance across varying context lengths. This diverse benchmark suite encompasses different long-context strategies including sparse attention, position encoding extensions, and memory-augmented architectures, ensuring robust evaluation of ERMAR’s retrieval-based approach against complementary methodologies.
3.1.5 Evaluation Metrics
We employ perplexity as the primary metric for language modeling performance, computed on the final 2048 tokens of each sequence to focus on long-range dependency modeling. For in-context learning tasks, we report accuracy on five natural language understanding benchmarks: SST-2, MR, Subj, SST-5, and MPQA, evaluated in both 4-shot and 20-shot settings. Memory efficiency is assessed through peak GPU memory usage and tokens processed per second, while computational overhead is measured via inference latency across different context lengths.
PG19 Proof-pile Wikitext-103
Model 1k 2k 4k 16k 1k 2k 4k 16k 1k 2k 4k 16k
7B Model
YARN-128k-7b 7.22 7.47 7.17 - 3.03 3.29 2.98 - 5.71 6.11 5.71 -
LongLoRA-7B-32k 9.76 9.71 10.37 7.62 3.68 3.35 3.23 2.60 7.99 7.83 8.39 5.47
LLaMA-2-7B 10.82 10.06 8.92 - 3.24 3.40 2.72 - 10.82 6.49 5.66 -
3B Model
Phi3-128k 11.31 9.90 9.66 -/9.65 4.25 3.11 2.77 -/3.08 7.54 7.22 7.01 -/7.20
OpenLLaMA-3B 11.60 9.77 $>$ $10^{3}$ 2.96 2.70 $>$ $10^{3}$ 10.57 8.08 $>$ $10^{3}$
LongLLaMA-3B* 10.59 10.02 $>$ $10^{3}$ 3.55 3.15 $>$ $10^{3}$ 8.88 8.07 $>$ $10^{3}$
LongLLaMA-3B 10.59 10.25 9.87 - 3.55 3.22 2.94 - 10.69 8.33 7.84 -
MemLong-3B* 10.66 10.09 $>$ $10^{3}$ 3.58 3.18 $>$ $10^{3}$ 8.72 7.93 $>$ $10^{3}$
w/ 4K MemLong 10.54 9.95 9.89 9.64 3.53 3.16 3.15 2.99 8.53 7.92 7.87 7.99
w/ 4K ERMAR 10.32 9.75 9.78 9.81 3.24 2.98 3.03 3.18 7.92 7.41 7.34 7.08
Table 1: Perplexity comparison of 7B and 3B models across PG19, Proof-pile, and WikiText-103, using a sliding window evaluation. "-" denotes Out of Memory (OOM) errors, and "x/y" indicates results from single/dual GPU setups. Memory-augmented models are tested with varying capacities. All runs use a single 3090 24GB GPU.
3.2 Results and Discussion
3.2.1 Long-Context Language Modeling
Following the experimental strategy adopted in Liu et al. (2024), Table 1 presents the mean perplexity scores of our model across different sequence lengths and datasets, demonstrating its effectiveness in long-context modeling. Evaluation was performed on test splits of three datasets: WikiText-103 Merity et al. (2016) (4,358 rows), PG-19 Rae et al. (2019) (100 rows), and Proof-Pile Azerbayev et al. (2023) (46.3k rows).
Among 7B models, YARN-128k-7B excels in shorter contexts, while LongLoRA-7B-32k scales effectively to 16k-token sequences, though with some performance degradation. This highlights the trade-off between performance and scalability, guiding model selection based on use-case needs.
The 3B models demonstrate ERMAR’s significant advantages in long-context tasks. While OpenLLaMA-3B struggles beyond 4k tokens, and Phi3-128k shows more consistent performance, ERMAR sets new performance benchmarks. ERMAR outperforms MemLong and larger 7B models in several configurations, including achieving a remarkable 2.98 perplexity on Proof-pile, surpassing MemLong’s 3.16. Despite slight underperformance in specific Proof-pile configurations, ERMAR’s enhanced memory retrieval mechanism proves more effective overall. Additionally, ERMAR maintains stable performance up to 16k tokens, with only a 3.2% degradation in perplexity from 4k to 16k tokens, demonstrating its superior scalability in long-context modelling.
3.2.2 Scalability to Extended Contexts
ERMAR’s scalability was evaluated at 32k tokens (Table 2). ERMAR demonstrates consistent performance advantages across multiple datasets: achieving a perplexity of 9.765 versus 9.858 on PG19 (0.90% improvement) and 7.880 versus 7.938 on WikiText-103 (0.06% improvement). Both methods achieve identical performance on Proof-pile with a perplexity of 3.063. This highlights ERMAR’s robustness for ultra-long contexts.
Dataset ERMAR MemLong Difference
PG19 9.765 9.858 0.90%
WikiText-103 7.880 7.938 0.06%
Proof-pile 3.063 3.063 0
Table 2: Perplexity at 32k context length, evaluated on NVIDIA L40S GPU.
3.2.3 In-Context Learning Performance
The results in Table 3 show ERMAR’s strong performance across five natural language understanding tasks in both 4-shot and 20-shot settings.
In the 4-shot setting, ERMAR achieves state-of-the-art results across all tasks, outperforming OpenLLaMA and other memory-augmented models. It excels even in challenging tasks like SST-5 and MPQA, maintaining high performance with limited examples. Its stability across different memory configurations highlights its robustness in low-resource scenarios.
ERMAR continues to excel in the 20-shot scenario, achieving top results in tasks like MPQA and Subj, and setting a new benchmark for SST-5. While it lags behind MemLong in a few tasks, ERMAR outperforms it overall, showcasing its scalability with increased examples.
ERMAR consistently performs well across varying context lengths, effectively leveraging memory augmentation. Its ability to scale with more examples and handle both short and long-range dependencies makes it a strong candidate for general-purpose language modelling, advancing the state-of-the-art in language understanding tasks.
[…]
OpenLLaMA 4,N/A 90.7 84.0 58.2 41.0 70.5 68.9
w./ Rag 4,4 90.9 90.5 61.6 39.2 63.2 69.1
LongLLaMA 4,4 90.4 83.9 64.3 40.0 64.2 68.6
MemLong 4,4 91.5 84.5 61.5 41.4 70.2 69.8
ERMAR 4,4 93.6 90.8 65.3 45.8 85.2 76.14
LongLLaMA 4,18 91.4 87.1 59.1 41.0 64.5 68.7
MemLong 4,18 91.0 89.6 61.7 43.5 69.4 71.0
ERMAR 4,18 93.6 90.8 65.3 45.9 85.2 76.16
OpenLLaMA 20,N/A 93.6 91.2 55.4 38.2 66.4 69.0
w./ Rag 20,18 92.2 91.3 75.8 39.8 57.6 71.3
LongLLaMA 20,18 94.1 90.8 64.2 41.4 72.1 72.7
MemLong 20,18 93.5 93.8 65.8 43.3 70.6 73.4
ERMAR 20,18 94.7 91.7 82.8 47 86.5 80.54
Table 3: 4-shot and 20-shot ICL accuracy [%] on 5 NLU tasks (SST-2, MR, Subj, SST-5, MPQA). We compare OpenLLaMA, LongLLaMA, MemLong, and ERMAR. Note: In-C = In-Context, In-M = In-Memory.
3.3 Memory Efficiency Analysis
Table 4 compares peak and reserved memory usage for ERMAR and MemLong across various context lengths, evaluated on an NVIDIA L40S (44.4GB) GPU.
Context Model Peak Mem Reserved Mem Mem/Token
Length (GB) (GB) (MB)
1024 ERMAR 7.97 8.16 7.97
MemLong 8.08 8.49 8.08
2048 ERMAR 8.45 8.71 4.22
MemLong 8.67 9.38 4.33
4096 ERMAR 9.42 9.87 2.35
MemLong 9.72 10.58 2.43
16384 ERMAR 15.20 16.61 0.95
MemLong 15.60 23.77 0.97
32768 ERMAR 22.87 25.56 0.71
MemLong 23.27 26.05 0.72
Table 4: Memory efficiency comparison of ERMAR and MemLong across context lengths on Wikitext data. Mem/Token is calculated as Peak Mem divided by context length.
ERMAR consistently uses less memory per token, particularly at longer contexts (16k and 32k), demonstrating its efficiency in memory management. The reduced reserved memory (e.g., 16.61 GB vs. 23.77 GB at 16k) underscores ERMAR’s optimized dynamic memory management, mitigating the information overload noted in the main text.
3.4 Latency and Throughput Analysis
Table 5 provides latency and throughput comparisons for ERMAR and MemLong on the PG-19 dataset, evaluated on an NVIDIA L40S (44.4GB) GPU.
Latency Throughput
Context Model Latency /Token (tokens/
(ms) sec)
1024 ERMAR 207.97 $\pm$ 53.01 0.203 5135
MemLong 190.68 $\pm$ 154.55 0.186 6803
2048 ERMAR 423.79 $\pm$ 52.56 0.206 4896
MemLong 323.49 $\pm$ 204.89 0.157 7446
4096 ERMAR 1358.36 $\pm$ 55.58 0.331 3020
MemLong 1184.16 $\pm$ 170.56 0.289 3511
16384 ERMAR 6862.98 $\pm$ 45.65 0.418 2387
MemLong 6496.00 $\pm$ 213.56 0.396 2524
32768 ERMAR 13679.03 $\pm$ 58.76 0.417 2395
MemLong 13449.90 $\pm$ 56.22 0.410 2436
Table 5: Latency and throughput comparison on PG-19 dataset. Latency is reported with standard deviation, and Throughput is in tokens per second.
ERMAR exhibits higher latency (9–35% at 1k–4k, narrowing to 1.7–5.5% at 16k–32k) but maintains competitive throughput ( 2.4k tokens/sec at long contexts). Notably, ERMAR’s lower latency variance (±45–59ms vs. MemLong’s ±154–214ms) indicates greater stability, aligning with its robustness in long-context tasks.
3.4.1 Single-Model Performance Scaling
To complement the comparative analysis, we evaluated ERMAR’s standalone performance characteristics across different sequence lengths on WikiText-103 with 16K memory capacity.
Seq Perplexity Memory Throughput Latency
Length /Token (tokens /Token
(GB) /sec) (ms)
1K 8.42 7.13 3125 0.32
2K 7.61 3.81 2904 0.35
4K 7.62 2.14 2109 0.47
8K 7.76 1.31 1836 0.54
16K 7.80 0.90 1727 0.58
Table 6: ERMAR’s standalone performance scaling on WikiText-103, showing memory efficiency and throughput characteristics across sequence lengths.
These results demonstrate ERMAR’s excellent scalability characteristics, with an 8-fold improvement in memory efficiency (7.13 → 0.90 MB/token) while maintaining stable perplexity performance. The throughput values on WikiText-103 are higher than the comparative PG-19 results, likely due to dataset-specific processing characteristics and the different evaluation methodologies used.
3.4.2 Ablation Studies
Ablation Study on Embedders: Following the comprehensive evaluation of BGE and LLM embedders, we expand our analysis to understand the nuanced impact of embedding architectures on ERMAR’s performance across different tasks and configurations.
We conducted extensive ablation studies comparing BGE and LLM embedders across SST-2, Subj, SST-5, MPQA, and MR tasks under various configurations, examining Flash vs. Eager attention, different context lengths (0 and 2048), and varying in-context/in-memory demonstration settings. The results are shown in Tables 7 and 8.
Flash Context In- In- SST-2 Subj SST-5 MPQA MR Average
Attention Length Context Memory
Flash 0 4 4 88.07 57.30 41.42 80.04 83.84 70.13
4 18 88.07 57.30 41.42 80.33 83.86 70.20
20 18 93.92 50.35 47.05 73.63 91.50 71.69
Eager 0 4 4 88.19 51.55 41.42 79.92 83.89 69.59
4 18 88.19 51.55 41.51 80.29 83.92 69.89
20 18 94.04 82.85 46.96 73.63 91.54 77.80
Flash 2048 4 4 88.07 51.55 41.42 80.04 83.84 68.98
4 18 88.07 51.25 41.42 80.33 83.86 68.99
20 18 93.92 82.75 47.05 73.63 91.50 77.77
Eager 2048 4 4 88.19 51.55 41.51 79.86 83.89 68.96
4 18 88.19 51.55 41.51 79.60 83.92 68.95
20 18 94.04 82.85 46.96 73.63 91.34 77.76
Table 7: Performance comparison of BGE embedder across configurations on SST-2, Subj, SST-5, MPQA, and MR tasks. Accuracy is reported in percentage.
Flash Context In- In- SST-2 Subj SST-5 MPQA MR Average
Attention Length Context Memory
Flash 0 4 4 88.07 57.30 41.42 80.04 83.84 70.13
4 18 88.07 57.30 41.42 80.33 83.86 70.20
20 18 93.92 50.35 47.05 73.63 91.54 71.70
Eager 0 4 4 88.19 57.30 41.51 79.86 83.89 70.15
4 18 88.19 57.30 41.51 79.60 83.92 70.10
20 18 94.04 50.35 46.96 73.63 91.54 71.70
Flash 2048 4 4 88.07 51.55 41.42 80.04 83.84 68.98
4 18 88.07 51.55 41.42 80.33 83.86 69.05
20 18 93.92 82.85 47.05 73.63 91.50 77.79
Eager 2048 4 4 88.19 51.25 41.51 79.92 83.89 68.95
4 18 88.19 51.25 41.51 80.29 83.92 69.03
20 18 94.04 82.75 46.96 73.63 91.34 77.74
Table 8: Performance comparison of LLM embedder across configurations on SST-2, Subj, SST-5, MPQA, and MR tasks. Accuracy is reported in percentage.
The ablation study reveals that both BGE and LLM embedders perform similarly, with the BGE embedder slightly outperforming LLM in the 20-shot, 18 in-memory setting (e.g., 77.80 vs. 71.70 for Eager attention at 0 context length). The performance difference is most pronounced in the Subj task, where BGE achieves up to 82.85% accuracy compared to LLM’s 50.35% in some configurations. This suggests that BGE may better capture semantic nuances in certain tasks, though both embedders maintain comparable performance overall. Further investigation into embedder-specific optimizations could enhance ERMAR’s performance, particularly for tasks requiring fine-grained contextual understanding.
Ablation Study on Relevance and Reranking:
We conducted detailed ablation studies on the core Relevance+Re-ranking mechanism, evaluating its impact on perplexity across varying context lengths for WikiText-103 and PG19 datasets. The results are summarized in Table 9.
Dataset Context Without With
Length Relevance+Reranker Relevance+Reranker
Perplexity Perplexity
Wiki 1K 8.841 7.919
2K 7.984 7.410
4K 7.438 7.437
16K 7.267 7.082
32K 7.938 8.008
PG19 1K 11.451 10.322
2K 10.412 9.746
4K 9.932 9.780
16K 9.910 9.809
32K 9.858 9.765
Table 9: Ablation study showing the impact of the Relevance+Reranking mechanism on perplexity for the WikiText-103 (Wiki) and PG19 datasets across different context lengths.
Relevance Mechanism Effectiveness: The relevance scoring mechanism shows strongest benefits at shorter context lengths: 10.4% improvement at 1K tokens and 7.2% at 2K tokens for WikiText-103. Benefits diminish at 4K tokens and become slightly negative at 32K tokens, suggesting ranking overhead outweighs advantages when memory capacity is sufficient. PG19 shows more consistent improvements (1.0%-9.9%) across all lengths, indicating narrative text benefits more from semantic ranking as story elements can be referenced non-sequentially.
4 Conclusion
We presented a novel ERMAR framework that enhances long-context modelling through relevance scoring and adaptive memory retrieval. ERMAR outperforms baseline models, including OpenLLaMA, LongLLaMA, and MemLong, achieving superior perplexity and in-context learning performance on WikiText-103, PG19, and Proof-pile. Future work will focus on optimizing ERMAR for specialized datasets and expanding its applicability to complex reasoning tasks.
5 Limitations
While ERMAR improves retrieval efficiency and context retention, it has limitations. Its reliance on ranked memory structures increases computational overhead compared to standard LLMs, particularly for large-scale retrieval as discussed in A.5 and A.6 in appendix sections. Additionally, performance variations across different task domains indicate a need for further tuning. The framework’s effectiveness in real-world, noisy environments also requires further validation.
Appendix
This appendix provides supplementary details for the Enhanced Ranked Memory Augmented Retrieval (ERMAR) framework, including dataset preprocessing, experimental hyperparameters and a glossary of terms
A.1 Dataset Preprocessing
The SlimPajama dataset, used for fine-tuning ERMAR, underwent several preprocessing steps to ensure compatibility with long-context tasks. The dataset was tokenized using the OpenLLaMA tokenizer, with a maximum sequence length of 32768 tokens. Duplicate sequences were removed using a hash-based deduplication algorithm, reducing the dataset to 84.7K unique training rows. To handle variable context lengths, we applied a sliding window approach with a stride of 512 tokens, ensuring that the model could process contexts ranging from 1024 to 32768 tokens. Special tokens were added to denote document boundaries, and padding was applied to align sequences to the nearest multiple of 128 tokens for efficient batch processing.
Training Configuration
ERMAR was trained using a two-stage fine-tuning approach with the following configuration:
Model Parameters:
  • Base model: OpenLLaMA-3B-v2 with LoRA adaptations
  • Memory layer: Layer 13 for historical context storage
  • Retrieval attention layers: [14, 18, 22, 26]
  • Memory capacity: 32,768 key-value pairs
  • Memory group size: 128 tokens per memory group
  • Retrieval group size: 8 (TopK retrieval)
  • Gate mechanism: Disabled (use_gate=False)
Training Hyperparameters:
  • Learning rate: $5\times 10^{-5}$ with 1,000 warmup steps
  • Weight decay: $1\times 10^{-4}$
  • Batch size: 1 per device (with gradient accumulation)
  • Sequence length: 1,024 tokens
  • Last context length: 1,024 tokens
  • Training epochs: 1 epoch on SlimPajama 0.5B subset
  • Training mode: LoRA-freeze (partial parameter updates)
LoRA Configuration:
  • Target modules: q_proj, k_proj, v_proj, o_proj
  • Trainable parameters: Layer normalization and embeddings
  • Frozen layers: Layers 0-13 (up to memory layer)
  • Position encoding: Zero position type for extended contexts
Embedder Setup:
  • Embedder: BAAI/bge-m3 (BGE embedder)
  • Embedding dimension: 1,024
  • GPU-based similarity search for efficient retrieval
Hardware and Infrastructure:
  • Primary training: Single NVIDIA 3090 24GB GPU
  • Extended context (32k): NVIDIA L40S 44.4GB GPU
  • Distributed training: ZeRO-2 optimization with Accelerate
  • Memory optimization: Sequential batching and continual fine-tuning
A.3 Glossary of Terms
To aid understanding, we provide definitions for key terms used in the ERMAR framework:
  • Relevance Score ($\alpha$): A normalized score computed via softmax over the dot product of query and key embeddings, representing the contextual importance of a memory entry (see Equation 1).
  • Key-Value Pair: A tuple $(K_{j},V_{j})$ storing contextual information, where $K_{j}$ is the key embedding and $V_{j}$ is the corresponding value embedding in the memory bank.
  • RSAR (Relevance Scoring with Adaptive Retrieval): The mechanism that dynamically ranks key-value pairs based on their relevance to the query, incorporating historical usage patterns.
  • Memory Bank: A non-trainable storage of pre-computed key-value embeddings, used to retain historical context without recomputation.
  • TopK Retrieval: The process of selecting the top-$K$ most relevant memory entries based on their relevance scores for use in the current context.
A.4 Fine-Grained Context Length Analysis
Figure 3: ERMAR perplexity performance across fine-grained context lengths for PG-19 and WikiText-103 datasets. The analysis reveals dataset-specific scaling patterns and validates performance stability across extended contexts.
Figure 3 presents detailed perplexity measurements across incremental context lengths from 1000 to 16000 tokens, providing fine-grained insights into ERMAR’s scaling behavior. The fine-grained analysis reveals that WikiText-103 achieves optimal performance around 13K-16K tokens (perplexity  7.36), while PG-19 maintains consistent performance (9.74-9.97) across all context lengths. This validates ERMAR’s robustness and suggests that factual content benefits more from extended context than narrative text.